# ProveRAG: Provenance-Driven Vulnerability Analysis with Automated Retrieval-Augmented LLMs

### v2

@@ 5.5 Improvement w/ Structured Context
Furthermore, to analyze the contribution of each URL and the Aqua database in generating summaries, Table 3 illustrates how many URLs the Retr. LLM used to generate summaries. These summaries were further incorporated into the Gen. LLM’s context to provide the exploitation and mitigation responses. For exploitation, it is shown that NVD and Aqua play a significant role, as they contain substantial exploitation-related information. Similarly, for mitigation, CWE and Aqua are the primary sources due to the dedicated section on mitigation strategies. Notably, the count for Others reaches 1130, as multiple hyperlinks may correspond to a single CVE entry. This demonstrates Aqua’s substantial contribution to both exploitation and mitigation responses, positively influencing ProveRAG’s overall performance (as shown in the bottom row of Table 2).
Figure 5: ProveRAG-Aqua attribution results
To further evaluate Aqua’s impact on the Gen. LLM’s responses, Figure 5 presents the attribution of each summary in ProveRAG-Aqua. The results reveal that, for the mitigation prompt, both Aqua and CWE exhibit a wide distribution of attention scores, with Aqua having a higher median around 0.7 compared to CWE’s median of 0.43. For the exploitation prompt, both Aqua and NVD display a broad range of attention scores, mostly positive, with Aqua again leading with a median near 0.7, while NVD’s median is around 0.3. This shows that when Aqua is included, it is the dominant source for the Gen. LLM for both exploitation and mitigation questions, with some contribution from NVD for exploitation and from CWE for mitigation strategies. Next, we will assess the performance of ProveRAG on another LLM.
5.6 ProveRAG’s Performance on Other LLMs (Generalizability)
@@
To evaluate whether the proposed ProveRAG framework is generalizable across different models, we compare the gpt-4o-mini results with Llama-3.1-8B (both models have the same context window and knowledge cutoff before the year 2024). First, based on qualitative evaluation, we noticed that gpt-4o-mini model is significantly better in following instructions to generate the responses. Note that the size of gpt-4o-mini model is not disclosed, but we argue that it is a much richer model in following instructions and generating in the desired format. On the other hand, Llama-3.1-8B model tend to not pay as much attention on the guidelines in generating the outputs, and many times failing to output the value, rationale, and provenance attributes in the right format (such as providing multiple TP/FP/FNs for a single response and/or not showing evidence for provenance). Therefore, we added the following to the Evaluation Prompt of ProveRAG for the Llama-3.1-8B model to alleviate this issue:
Output in the following format: value: TP or FP or FN (ONLY one value for the entire response) rationale: [the rationale segment for the selected value] provenance: - response: [extract the relevant segment from the response segment]’ - context: [extract the relevant segment from the context segment]’
This indicates the limitation for LLMs with different sizes and capabilities, as they may struggle to follow all the instructions when tasked with a lengthy or complex set of instructions. However, through careful prompt engineering, this challenge can be substantially mitigated.
@@
Another qualitative observation from Llama-3.1-8B is that the guardrails are more strict for this model. This was not the case for the gpt-4o-mini model as it never refused to generate the response for the Generation Prompt. However, Llama-3.1-8B generated the following sentence for all the CVEs for the prompt-only experiment:
@@
This was mostly an issue for the prompt-only experiment, however, for ProveRAG (i.e., with retrieved and summarized information injected into the prompt), the Llama-3.1-8B only refused to generate the response for around 40 of the CVEs. Therefore, we did not include those 40 for provenance (as provenance requires a response to self-critique and provide evidence). This shows that different LLMs may have varying sets of guardrails, but with careful prompting, they can still benefit from the ProveRAG framework. Table 4 shows the results obtained. Note that the #GR indicates the number of guardrail responses, which means that the LLM refused to provide a response for them.
Table 4: Performance of ProveRAG for GPT-4o-mini vs. Llama3.1-8B Model
@@
Based on the results, it is evident that ProveRAG performs significantly better with both the gpt-4o-mini and Llama-3.1-8B models compared to the baselines, with gpt-4o-mini showing particularly superior performance. This shows ProveRAG’s effectiveness in vulnerability analysis on another LLM, showcasing its ability to identify CVEs and conduct self-critique. Moreover, our findings indicated that Llama-3.1-8B tend to “forget” to adhere to all given instructions. Therefore, we recommend a more careful prompt engineering to enhance performance, which allows the model to better follow instructions and bypass overly rigid guardrails through the proposed ProveRAG framework. While the choice of model is flexible, we suggest using models with strong instruction-following capabilities for optimal outcomes.
@@
Therefore, these findings support the generalizability of the proposed ProveRAG framework, emphasizing that with careful prompt engineering—taking into account the unique guardrails and instruction-following capabilities of each model—ProveRAG can effectively leverage different LLMs. To optimize the effectiveness of provenance, we recommend using models that excel at capturing complex instructions and understanding intricate token dependencies. In particular, prompts must be tailored to each model’s specific characteristics to fully leverage the benefits of ProveRAG. Next, we will show some specific examples on how the proposed ProveRAG framework works.
5.7 Specific Examples
@@
In this section, we show how the evaluation module of ProveRAG works (RQ3). The first three examples present gpt-4o-mini responses with varying TP, FP, and FN predictions. Following these, we provide a comparative analysis of a specific CVE between the gpt-4o-mini and Llama-3.1-8B models. First, here is one example of gpt-4o-mini that shows a TP response:
@@
CVE-2024-0244 — Out-of-bounds Write –> GPT Mitigation Strategies value: “TP”, rationale: The response accurately reflects the mitigation strategies for CVE-2024-0244 as outlined in the context. It mentions updating the firmware, avoiding direct internet connections, using secure private networks, and employing various programming practices to prevent buffer overflows, all of which are supported by the context provided. provenance: response: To mitigate this vulnerability, it is recommended to update the firmware of the affected devices to versions beyond 03.07, as advised by Canon. Additionally, users should avoid direct internet connections and instead use a secure private network with a firewall. Other strategies include using programming languages that prevent buffer overflows, employing vetted libraries, implementing automatic buffer overflow detection mechanisms, and adhering to strict memory management practices. context: …we advise that our customers install the latest firmware available for the affected models… We also recommend that customers set a private IP address for their products and create a network environment with a firewall or Wired/Wi-Fi router that can restrict network access. Potential Mitigations… Use a language that does not allow this weakness to occur or provides constructs that make this weakness easier to avoid… Use a vetted library or framework that does not allow this weakness to occur or provides constructs that make this weakness easier to avoid… Use automatic buffer overflow detection mechanisms that are offered by certain compilers or compiler extensions. ——————– Retrieved URLs: nvd.nist.gov, cwe.mitre.org, avd.aquasec.com, psirt.canon/advisory-information/cp2024-001, canon.jp/support/support-info/240205vulnerability-response, canon-europe.com/support/product-security-latest-news ——————– Metrics: Cosine Similarity: 0.7607, ROUGE-L: 0.2299
@@
The above example shows that ProveRAG-Aqua successfully provides an accurate (TP) response, with a rationale and substantiated provenance. Interestingly, both the response and a significant portion of the evidence were derived from the hyperlinks, specifically “canon.jp”. This underscores the critical importance of integrating a wide array of URLs to ensure a comprehensive coverage that enhances the system’s capability to provide well-supported evidence. The provided metrics correspond to the values shown in Figure 3. Here is an hallucination (FP) example of gpt-4o-mini:
@@
CVE-2024-0338 — Classic Buffer Overflow –> GPT Exploitation Information value: “FP”, rationale: The response inaccurately describes the exploitation of CVE-2024-0338. It mentions a general process of exploitation involving a flaw in the system, but it fails to specify that this particular vulnerability is a buffer overflow in XAMPP, which allows arbitrary code execution through a long file debug argument. The response does not accurately reflect the specific nature of the vulnerability as described in the context. provenance: response: CVE-2024-0338 is a vulnerability that allows attackers to exploit a flaw in the system, potentially leading to unauthorized access or data manipulation. context: A buffer overflow vulnerability has been found in XAMPP affecting version 8.2.4 and earlier. An attacker could execute arbitrary code through a long file debug argument that controls the Structured Exception Handler (SEH). ——————– Metrics: Cosine Similarity: 0.4075, ROUGE-L: 0.1695
@@
CVE-2024-0925 — Stack Buffer Overflow –> GPT Mitigation Strategies value: “FN”, rationale: The response omits specific details about the potential mitigations that are present in the context. While it mentions general strategies like buffer overflow detection mechanisms and ASLR, it does not include the specific examples and additional context provided in the context, such as the specific compiler flags and techniques like StackGuard, ProPolice, and D3-SFCV. provenance: response: Potential mitigations include using automatic buffer overflow detection mechanisms, employing Address Space Layout Randomization (ASLR), and compiling software with features that randomize memory addresses to prevent predictable exploit paths. context: Use automatic buffer overflow detection mechanisms that are offered by certain compilers or compiler extensions. Examples include: the Microsoft Visual Studio /GS flag, Fedora/Red Hat FORTIFY_SOURCE GCC flag, StackGuard, and ProPolice, which provide various mechanisms including canary-based detection and range/index checking. ——————– Retrieved URLs: nvd.nist.gov, cwe.mitre.org, avd.aquasec.com, https://github.com/yaoyue123/iot/blob/main/Tenda/AC10U/ ——————– Metrics: Cosine Similarity: 0.7038, ROUGE-L: 0.1892
@@
In this case, the response generated by ProveRAG-Aqua is ‘FN’ since the LLM omitted some detailed information regarding the mitigation strategies. Note that the response is actually not bad; it just lacks some details, and the metrics reflect that it is actually close to the context in the Retrieved URLs. This underscores the closeness between TP and FN values shown in Figure 3.
Now, to see the performance of ProveRAG on other LLM (and provenance quality), here are the responses of CVE-2024-0267 for both Llama3.1-8B and gpt-4o-mini:
@@
CVE-2024-0267 — SQL Injection –> Llama Exploitation Information value: “FN”, rationale: This CVE (2024-0267) is a known vulnerability, and it has been disclosed to the public. The associated identifier of this vulnerability is VDB-249823. provenance: response: A vulnerability classified as critical was found in Kashipara Hospital Management System up to 1.0. Affected by this vulnerability is an unknown functionality of the file login.php of the component Parameter Handler. context: New CVE Received from VulDB ——————– Retrieved URLs: nvd.nist.gov, vuldb.com/?ctiid.249823 ——————– Metrics: Cosine Similarity: 0.3070, ROUGE-L: 0.0000
@@
CVE-2024-0267 — SQL Injection –> GPT Exploitation Information value: “TP”, rationale: The response accurately describes the critical vulnerability CVE-2024-0267 in the Kashipara Hospital Management System, specifically detailing the SQL injection through the ’email’ and ’password’ parameters in the login.php file. It correctly states that an attacker can exploit this vulnerability to gain unauthorized access to the system’s database, which aligns with the context provided. provenance: response: CVE-2024-0267 is a critical vulnerability in the Kashipara Hospital Management System that allows for SQL injection through the manipulation of the ’email’ and ’password’ parameters in the login.php file. An attacker can exploit this vulnerability by injecting malicious SQL code into these parameters, which can lead to unauthorized access to the system’s database. context: A vulnerability classified as critical was found in Kashipara Hospital Management System up to 1.0. This vulnerability is known as CVE-2024-0267. The manipulation of the argument email/password leads to sql injection. The attack can be launched remotely. ——————– Retrieved URLs: nvd.nist.gov, cwe.mitre.org, vuldb.com/?id.249823 ——————– Metrics: Cosine Similarity: 0.8785, ROUGE-L: 0.3265
@@
For CVE-2024-0267 with Llama-3.1-8B, the prediction is classified as a “FN” for exploitation information. The rationale and provenance validate the selected value in ProveRAG. As can be seen in the context, the Llama model failed to properly extract the retrieved information. Low Rouge-L and Cosine Similarity scores indicate the difference between the generated response and the supporting evidence. This shows how LLMs with weak instruction-following abilities will perform for provenance, and hence, indicating the need for either a more powerful model or a careful engineered prompt tailored to this model’s specific characteristics. In contrast, the gpt-4o-mini model accurately identified the prediction as “TP” by providing evidence from the Retrieved URLs, in which have strong overlaps with the model’s response with high Cosine Similarity and Rouge-L scores. These metrics correspond to the values presented in Figure 3. This comparison demonstrates how different models may struggle to effectively utilize retrieved context, emphasizing the role of model selection and/or a tailored, engineered prompt for a specific model to maximize the benefits of ProveRAG in optimizing performance and reliability in vulnerability analysis.
6 Conclusion

### v3

@@ V-D Improvement w/ Structured Context
Furthermore, to analyze the contribution of each URL and the Aqua database in generating summaries, Table III illustrates how many URLs the Retr. LLM used to generate summaries. These summaries were further incorporated into the Gen. LLM’s context to provide the exploitation and mitigation responses. For exploitation, it is shown that NVD and Aqua play a significant role, as they contain substantial exploitation-related information. Similarly, for mitigation, CWE and Aqua are the primary sources due to the dedicated section on mitigation strategies. Notably, the count for Others reaches 1130, as multiple hyperlinks may correspond to a single CVE entry. This demonstrates Aqua’s substantial contribution to both exploitation and mitigation responses, positively influencing ProveRAG’s overall performance (as shown in Table II).
V-E ProveRAG’s Generalizability on Another LLM
To evaluate whether the proposed ProveRAG framework is generalizable across different models, we compare the gpt-4o-mini results with Llama-3.1-8B (both models have the same context window and knowledge cutoff before the year 2024). It is worth noting that we conducted our experiments with Llama-3.1-8B on a system equipped with an Intel Xeon W-2145 CPU (8 cores, 3.7GHz), an NVIDIA GeForce RTX 3060 GPU with 12 GB memory, and 32 GB of RAM.
@@ V-E ProveRAG’s Generalizability on Another LLM
First, based on qualitative evaluation, we noticed that gpt-4o-mini model is significantly better in following instructions to generate the responses. Note that the size of gpt-4o-mini model is not disclosed, but we argue that it is a much richer model in following instructions and generating in the desired format. On the other hand, Llama-3.1-8B model tends to not pay as much attention on the guidelines in generating the outputs, and often fails to output the value, rationale, and provenance attributes in the right format (such as providing multiple TP/FP/FNs for a single response and/or not showing evidence for provenance). Therefore, we added the following to the Evaluation Prompt of ProveRAG for this model to alleviate this issue:
{promptbox}
Output in the following format:
value: TP or FP or FN (ONLY one value for the entire response)
rationale: [the rationale segment for the selected value]
provenance:
response: [extract the relevant segment from the response segment]
context: [extract the relevant segment from the context segment]
@@
This indicates the limitation for LLMs with different sizes and capabilities, as they may struggle to follow all the instructions when tasked with a lengthy or complex set of instructions. However, through careful prompt engineering, this challenge can be mitigated. Another observation from deploying ProveRAG with Llama-3.1-8B is that the guardrails are more strict. This was not the case for gpt-4o-mini as it never refused to generate the response for the Generation Prompt. On the other hand, during the prompt-only experiment, Llama-3.1-8B generated the following sentence for all the CVEs:
{promptbox}
@@
This was mainly an issue for the prompt-only experiment. With ProveRAG, where the retrieved and summarized information is injected into the prompt, Llama-3.1-8B only refused to generate the response for around 40 of the CVEs. We excluded these 40 cases from the provenance evaluation, as the self-critique mechanism requires a generated response from the sources to compute metrics (TP/FP/FN) and provide evidence. This shows that different LLMs may have varying sets of guardrails, but with careful prompting, they can still benefit from the ProveRAG framework.
Table IV shows the results obtained across the evaluated models. Note that #GR indicates the number of guardrailed responses, which means that the LLM refused to provide a response for them. Based on the results, it is evident that ProveRAG performs significantly better with both the gpt-4o-mini and Llama-3.1-8B models compared to the baselines, with gpt-4o-mini showing particularly superior performance. This shows ProveRAG’s effectiveness in vulnerability analysis on another LLM, showcasing its ability to identify CVEs and conduct self-critique.
TABLE IV: Performance of ProveRAG for GPT-4o-mini vs. Llama3.1-8B Model
@@
Moreover, our findings indicated that Llama-3.1-8B tend to “forget” to adhere to all given instructions. Therefore, we recommend a more careful prompt engineering to enhance performance, which allows the model to better follow instructions and bypass overly rigid guardrails through the proposed ProveRAG framework. While the choice of model is flexible, we suggest using models with strong instruction-following capabilities for optimal outcomes. Note that one limitation of our study is the computational budget required to test larger open-source models, and that is why we only used Llama-3.1-8B to compare against gpt-4o-mini.
@@
These findings support the generalizability of the proposed ProveRAG framework, emphasizing that with careful prompt engineering—taking into account the unique guardrails and instruction-following capabilities of each model—ProveRAG can effectively leverage different LLMs. To optimize the effectiveness of provenance, we recommend using models that excel at capturing complex instructions and understanding intricate token dependencies. In particular, prompts must be tailored to each model’s specific characteristics to fully leverage the benefits of ProveRAG. Next, we will show some specific examples on how the ProveRAG framework works.
V-F Specific Examples
@@
In this section, we show how the evaluation module of ProveRAG works (RQ3). The first three examples present gpt-4o-mini responses with TP, FP, and FN predictions. Following these, we provide a comparative analysis of a specific CVE between the gpt-4o-mini and Llama-3.1-8B models. First, here is one example of gpt-4o-mini that shows a TP response:
{mybox}
[frametitle=CVE-2024-0244 — Out-of-bounds Write –¿ GPT]
Mitigation Strategies
Value: “TP”,
Rationale: The response accurately reflects the mitigation strategies for CVE-2024-0244 as outlined in the context. It mentions updating the firmware, avoiding direct internet connections, using secure private networks, and employing various programming practices to prevent buffer overflows, all of which are supported by the context provided.
Provenance:
@@
response: To mitigate this vulnerability, it is recommended to update the firmware of the affected devices to versions beyond 03.07, as advised by Canon. Additionally, users should avoid direct internet connections and instead use a secure private network with a firewall. Other strategies include using programming languages that prevent buffer overflows, employing vetted libraries, implementing automatic buffer overflow detection mechanisms, and adhering to strict memory management practices.
context: …we advise that our customers install the latest firmware available for the affected models… We also recommend that customers set a private IP address for their products and create a network environment with a firewall or Wired/Wi-Fi router that can restrict network access. Potential Mitigations… Use a language that does not allow this weakness to occur or provides constructs that make this weakness easier to avoid… Use a vetted library or framework that does not allow this weakness to occur or provides constructs that make this weakness easier to avoid… Use automatic buffer overflow detection mechanisms that are offered by certain compilers or compiler extensions.
Retrieved URLs:
nvd.nist.gov, cwe.mitre.org, avd.aquasec.com,
psirt.canon/advisory-information/cp2024-001,
canon.jp/support/support-info/240205vulnerability-response,
canon-europe.com/support/product-security-latest-news
Metrics:
Cosine Similarity: 0.7607, ROUGE-L: 0.2299
@@
The above example shows that ProveRAG-Aqua successfully provides an accurate (TP) response, with a rationale and substantiated provenance. Interestingly, both the response and a significant portion of the evidence were derived from the hyperlinks, specifically “canon.jp”. This underscores the critical importance of integrating a wide array of URLs to ensure a comprehensive coverage that enhances the system’s capability to provide well-supported evidence. In practice, this translates into responses that not only achieve factual accuracy but also provide transparency and traceability through verifiable provenance, enabling end-users to independently confirm the validity of the recommendations. The provided metrics correspond to the values shown in Figure 3. Here is an hallucination (FP) example detected by gpt-4o-mini:
{mybox}
[frametitle=CVE-2024-0338 — Classic Buffer Overflow –¿ GPT]
Exploitation Information
Value: “FP”,
@@
Rationale: The response inaccurately describes the exploitation of CVE-2024-0338. It mentions a general process of exploitation involving a flaw in the system, but it fails to specify that this particular vulnerability is a buffer overflow in XAMPP, which allows arbitrary code execution through a long file debug argument. The response does not accurately reflect the specific nature of the vulnerability as described in the context.
Provenance:
response: CVE-2024-0338 is a vulnerability that allows attackers to exploit a flaw in the system, potentially leading to unauthorized access or data manipulation.
context: A buffer overflow vulnerability has been found in XAMPP affecting version 8.2.4 and earlier. An attacker could execute arbitrary code through a long file debug argument that controls the Structured Exception Handler (SEH).
Metrics:
Cosine Similarity: 0.4075, ROUGE-L: 0.1695
@@
{mybox}
[frametitle=CVE-2024-0925 — Stack Buffer Overflow –¿ GPT]
Mitigation Strategies
Value: “FN”,
Rationale: The response omits specific details about the potential mitigations that are present in the context. While it mentions general strategies like buffer overflow detection mechanisms and ASLR, it does not include the specific examples and additional context provided in the context, such as the specific compiler flags and techniques like StackGuard, ProPolice, and D3-SFCV.
Provenance:
response: Potential mitigations include using automatic buffer overflow detection mechanisms, employing Address Space Layout Randomization (ASLR), and compiling software with features that randomize memory addresses to prevent predictable exploit paths.
context: Use automatic buffer overflow detection mechanisms that are offered by certain compilers or compiler extensions. Examples include: the Microsoft Visual Studio /GS flag, Fedora/Red Hat FORTIFY_SOURCE GCC flag, StackGuard, and ProPolice, which provide various mechanisms including canary-based detection and range/index checking.
Retrieved URLs:
nvd.nist.gov, cwe.mitre.org, avd.aquasec.com,
https://github.com/yaoyue123/iot/blob/main/Tenda/AC10U/
Metrics:
Cosine Similarity: 0.7038, ROUGE-L: 0.1892
@@
In this case, the response generated by ProveRAG-Aqua is ‘FN’ since the LLM omitted some detailed information regarding the mitigation strategies. Note that the response is actually not bad; it just lacks some details, and the metrics reflect that it is actually close to the context in the Retrieved URLs. This explains the closeness between TP and FN values shown in Figure 3. Now, to see the performance of ProveRAG on another LLM, here are the responses of CVE-2024-0267 for both Llama3.1-8B and gpt-4o-mini:
{mybox}
[frametitle=CVE-2024-0267 — SQL Injection –¿ Llama]
Exploitation Information
Value: “FN”,
Rationale: This CVE (2024-0267) is a known vulnerability, and it has been disclosed to the public. The associated identifier of this vulnerability is VDB-249823.
Provenance:
@@
response: A vulnerability classified as critical was found in Kashipara Hospital Management System up to 1.0. Affected by this vulnerability is an unknown functionality of the file login.php of the component Parameter Handler.
context: New CVE Received from VulDB
Retrieved URLs:
nvd.nist.gov, vuldb.com/?ctiid.249823
Metrics:
Cosine Similarity: 0.3070, ROUGE-L: 0.0000
{mybox}
[frametitle=CVE-2024-0267 — SQL Injection –¿ GPT]
Exploitation Information
Value: “TP”,
@@
Rationale: The response accurately describes the critical vulnerability CVE-2024-0267 in the Kashipara Hospital Management System, specifically detailing the SQL injection through the ’email’ and ’password’ parameters in the login.php file. It correctly states that an attacker can exploit this vulnerability to gain unauthorized access to the system’s database, which aligns with the context provided.
Provenance:
response: CVE-2024-0267 is a critical vulnerability in the Kashipara Hospital Management System that allows for SQL injection through the manipulation of the ’email’ and ’password’ parameters in the login.php file. An attacker can exploit this vulnerability by injecting malicious SQL code into these parameters, which can lead to unauthorized access to the system’s database.
context: A vulnerability classified as critical was found in Kashipara Hospital Management System up to 1.0. This vulnerability is known as CVE-2024-0267. The manipulation of the argument email/password leads to sql injection. The attack can be launched remotely.
Retrieved URLs:
nvd.nist.gov, cwe.mitre.org, vuldb.com/?id.249823
Metrics:
Cosine Similarity: 0.8785, ROUGE-L: 0.3265
@@
For CVE-2024-0267 with Llama-3.1-8B, the prediction is classified as a “FN” for exploitation information. The rationale and provenance validate the selected value in ProveRAG. As can be seen in the context, the Llama model failed to properly extract the retrieved information. Low Rouge-L and Cosine Similarity scores indicate the difference between the generated response and the supporting evidence. This shows how LLMs with weak instruction-following abilities will perform for provenance, and hence, indicating the need for either a more powerful model or a careful engineered prompt tailored to this model’s specific characteristics. In contrast, the gpt-4o-mini model accurately identified the prediction as “TP” by providing evidence from the Retrieved URLs, in which have strong overlaps with the model’s response with high Cosine Similarity and Rouge-L scores. These metrics correspond to the values presented in Figure 3.
This comparison demonstrates that different models can struggle to effectively leverage extensive retrieved context, which can introduce reliability limitations in LLM-based self-evaluation. Large amounts of information may lead the model to overlook or misprioritize relevant evidence, causing biased assessments. This emphasizes the importance of careful model selection and tailored, engineered prompts for each specific model to maximize the benefits of ProveRAG in optimizing performance and reliability in vulnerability analysis.
VI Conclusion